Assume Hardy-Weinberg equilibrium. In a population with 1000 individuals, if genotype frequencies are equal to AA 50%, Aa 42%, and aa 8%, what is the probability of obtaining a "AA" Zygote?

  

7%

   

50%

   

71%

   

5%
Answer #1

50%

AA = 50% = 0.5

A = √0.5

AA zygote = √0.5 × √0.5 = 0.5

Percentage of AA = 0.5 × 100 = 50%

Please rate high.